OpenStudy (maniacpokemon43):

the volume of a cylinder is 841 cubic ft. Find the height, rounding answer to the nearest hundredth (if necessary) RADIUS =18 ft

OpenStudy (anonymous):

screw this website, you will not get ANY help here. Some freaking bullpellet, go to yahoo answers. It is MUCH better....

OpenStudy (maniacpokemon43):

@TheSmartOne @Preetha

OpenStudy (maniacpokemon43):

@Mehek14

OpenStudy (maniacpokemon43):

@dan815 @zepdrix

Mehek (mehek14):

volume = \(πr^2h\) you have \(841 = π*256*h\) \(π * 256 = 803.84 \) \(841 - 803.84 = 37.16\)

OpenStudy (maniacpokemon43):

WHERE DID U GET 256

Mehek (mehek14):

\(16^2 = 256\)